OpenStudy (anonymous):

Find a cubic function with the given zeros. -2, 5, -6 i get u put (x+2)+(x-5)....... im just not sure what to do after that

OpenStudy (anonymous):

3(x+1) is what i got

OpenStudy (abdullahm):

That's wrong.

OpenStudy (abdullahm):

If r is a zero of a function, then (x-r) is a factor of the function. So, if we have -2, 5, and -6 as the zeros of a function, then what are all of the factors of the function? :)

OpenStudy (serenity74):

I am sorry you don't add them you mutiply them.

OpenStudy (abdullahm):

Exactly, why the answer was wrong.

OpenStudy (abdullahm):

Exactly why the answer was wrong.*

OpenStudy (serenity74):

No, if you add them together you get 3(x+1)

OpenStudy (serenity74):

When you are actually solving for Cubic functions with given zeros you multply all of them together in order to get an equation.

OpenStudy (abdullahm):

There is no reason to add them whatsoever. We need to find a cubic function. That means the highest degree is 3.

OpenStudy (serenity74):

Ik that's why I reposted my explaniation....

OpenStudy (abdullahm):

@jon321 Do you understand?

OpenStudy (anonymous):

no idea what just happened. these are possible answers f(x) = x3 + 3x2 + 28x - 60 f(x) = x3 - 3x2 - 28x - 60 f(x) = x3 + 3x2 - 28x - 60 f(x) = x3 + 3x2 - 28x + 60

OpenStudy (anonymous):

i think its C

OpenStudy (abdullahm):

If r is a zero of a function, then (x-r) is a factor of the function. So, if we have -2, 5, and -6 as the zeros of a function, then what are all of the factors of the function?

OpenStudy (abdullahm):

@jon321 You're correct! C is the answer. :)

OpenStudy (serenity74):

That is correct but do you know why? Cause we are all here to teach you and help you gain the knowledge.

OpenStudy (wolf1728):

To get that equation you would multiply these 3 terms (x+2)*(x-5)*(x+6)